When 105.7 g of H2O at 22ºC is mixed with an unknown mass of H2O at a 58ºC, the final temperature of the resulting mixture is 45ºC. What was the mass of the second sample of water? The specific heat of H20 is 4.184 J/g.ºC.
Answer #1

Q= 4.184joules for 1 gram at 1 degrees Celsius

Than Q for 1 gram at 45 degrees Celsius = 45× 4.184 j/g

Specific heat = Q/m×(t1 - t2)

4.184 = Q/(105.7+X) (13)

105.7+X = Q / 4.184 ×13